OBSERVATION OF THE PURE ROTATIONAL SPECTRA OF THE H_{2}O-trans-HOCO COMPLEX

\underline{TAKAHIRO OYAMA}, MASAKAZU NAKAJIMA, YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo, 153-8902, Japan; and YOSHIHIRO SUMIYOSHI, Department of Chemistry and Chemical Biology, Gunma University, 4-2 Aramaki-machi, Maebashi City, Gunma, 371-8510 Japan.

Rotational spectra of the H_{2}O-trans-HOCO complex have been observed using a Fourier transform microwave (FTMW) spectrometer. The complex was produced in a supersonic jet by discharging a gas mixture of CO and H_{2}O diluted in Ar. The observed lines show that the ground state of the complex is split into two by the exchange symmetry of the two equivalent protons of H_{2}O. The molecular constants including the hyperfine coupling constants have been precisely determined for the two states. The Fermi contact constants of the two states are smaller than that of the trans-HOCO monomer. This result indicates that there is an induced effect for the spin density on the hydrogen nucleus of HOCO by the complex formation.